The length of a rectangle is twice its width.
If the perimeter of the rectangle is 54 m , find its area

Answer by TimothyLamb(4379)

L = 2w
p = 2L + 2w = 54
---
2L + 2w = 54
2(2w) + 2w = 54
4w + 2w = 54
6w = 54
w = 9
L = 2w = 18
---
a = Lw
a = 18 * 9
---
Answer:
a = 162 sq.m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php